OpenStudy (anonymous):

given real numbers a,b,c,d>1 prove ((a^2)/(c-1))+((b^2)/(d-1))+((c^2)/(e-1))+((d^2)/(a-1))+((e^2)/(b-1)) >=20

OpenStudy (anonymous):

\[\frac{ a^2 }{ c-1 } +\frac{b^2}{d-1}+\frac{c^2}{e-1}+\frac{d^2}{a-1}+\frac{e^2}{b-1}\ge20\]

OpenStudy (kc_kennylau):

you mean integers

OpenStudy (anonymous):

no all real numbers including integers

OpenStudy (anonymous):

help guys

OpenStudy (anonymous):

help guys

OpenStudy (anonymous):

do it for two variables using gradient, you get the minimum of the sum to be 8 and achieved at a=2, b=2 \[ \frac{a^2}{b-1}+\frac{b^2}{a-1}\geq 8 \] for a, b>1

OpenStudy (anonymous):

The general case is messy. Maybe someone else can find an easier method, What is your background in Math?

OpenStudy (anonymous):

i am just 11th standard student

OpenStudy (anonymous):

and what is this gradient thing?

OpenStudy (anonymous):

OK, so there must be a must be a method not using Calculus. I will do some more work.